# AI-enhanced collective intelligence

### v3

@@
Hao Cui School of Sociology, University College Dublin, Belfield, Dublin 4, Dublin, Ireland Geary Institute for Public Policy, University College Dublin, Belfield, Dublin 4, Dublin, Ireland Taha Yasseri School of Sociology, University College Dublin, Belfield, Dublin 4, Dublin, Ireland Geary Institute for Public Policy, University College Dublin, Belfield, Dublin 4, Dublin, Ireland
@@
The current societal challenges exceed the capacity of human individual or collective effort alone. As AI evolves, its role within human collectives is poised to vary from an assistive tool to a participatory member. Humans and AI possess complementary capabilities that, when synergized, can achieve a level of collective intelligence that surpasses the collective capabilities of either humans or AI in isolation. However, the interactions in human-AI systems are inherently complex, involving intricate processes and interdependencies. This narrative review incorporates perspectives from complex network science to conceptualize a multilayer representation of human-AI collective intelligence, comprising cognition, physical, and information layers. Within this multilayer network, humans and AI agents exhibit varying characteristics; humans differ in diversity from surface-level to deep-level attributes, while AI agents range in degrees of functionality and anthropomorphism. The interplay among these agents shapes the overall structure and dynamics of the system. We explore how agents’ diversity and interactions influence the system’s collective intelligence. Furthermore, we present an analysis of real-world instances of AI-enhanced collective intelligence. We conclude by addressing the potential challenges in AI-enhanced collective intelligence and offer perspectives on future developments in this field.
@@ Introduction
The definition of artificial intelligence has undergone numerous evolutions in the past semi-century. Artificial intelligence is the simulation of human intelligence processes by machines 23, especially computer systems. Previous research considers the classification of AI in distinct ways 24. The first classification categorizes AI based on its human-like cognitive abilities, such as “think” and “feel.” In another categorization, there are four primary AI types 25: reactive AI, limited memory AI, theory of mind AI, and self-aware AI. Limited-memory AI is the most common at the current time, ranging from virtual assistants to chatbots to self-driving vehicles. Such AI can learn from historical data to recognize patterns, generate new knowledge and understanding, and inform subsequent decisions.
Yet, another classification is a technology-oriented approach that categorizes AI into Artificial Narrow Intelligence (ANI), Artificial General Intelligence (AGI), and Artificial Superintelligence (ASI). ANI involves creating computer systems capable of performing specific tasks like human intelligence 26, but often surpass humans in efficiency and accuracy 27. ANI machines have a narrow range of capabilities and can represent all existing AI. AGI refers to machines that exhibit the human ability to learn, perceive, and understand a wide range of intellectual tasks 28, 29. Finally, ASI’s primary goal is to develop a machine with cognitive abilities higher than those of humans 24.
Existing AI can perform tasks including but not limited to complex calculations, language translation, facial recognition, and financial market prediction. Recent developments in generative AI, such as OpenAI’s text-to-video model Sora 30, exemplify AI’s potential in creative industries. As AI technology evolves, it is poised to bring groundbreaking advancements across various fields. Beyond content creation, as AI deepens the understanding of the physical world and develops its simulation ability at the perceptual and cognitive level, it is possible to develop more super-intelligent tools in various fields.
Current AI systems can process vast amounts of data on a scale far beyond human capabilities. However, many real-world challenges cannot yet be solved solely by AI. Currently, AI lacks the deep conceptual and emotional understanding humans possess about objects and experiences 31. AI cannot interpret human language’s nuances, including contextual, symbolic, and cultural meanings 32, 33. AI cannot be an ethical decision-maker because it lacks the human attributes of intentionality, care, and responsibility 31. Therefore, combining human insight with AI’s analytical power is crucial for addressing complex real-world challenges, leveraging the strengths to compensate for each other’s weaknesses.
@@
Even though the field of collective intelligence initially focused solely on groups of people, in recent years, it has gradually expanded to include AI as group members in a new framework referred to as “hybrid intelligence” 34. Proponents of the hybrid intelligence perspective stress that humans and AI can connect in ways that allow them to collectively act more intelligently and achieve unreachable goals by any individual entities alone 35. Researchers investigating this hybrid collective intelligence explore “how people and computers can be connected so that collectively they act more intelligently than any person, group, or computer has ever done before” 36.
Acknowledging the complementary capabilities of humans and AI as discussed in the previous section, researchers identify the need for developing socio-technological ensembles of humans and intelligent machines that possess the ability to collectively achieve superior results and continuously improve by learning from each other 37. Previous research provides evidence that teaming humans with AI has the potential to achieve complementary team performance (CTP), a level of performance that AI or humans cannot reach in isolation 38. For instance, a study demonstrates that humans can use contextual information to adjust the AI’s decision 39. Research on mixed teams composed of humans and AI shows hybrid teams could achieve higher performance metrics, such as team situational awareness and score, than all-human teams 38.
@@
For example, machine learning and automation can increase the efficiency and scalability of collective intelligence in Citizen Science projects, where citizens volunteer to help scientists tag and classify their large-scale datasets 40. Members of a crowd with different interests and areas of competence can be matched to other tasks by recommendation systems 41. Generative algorithms can extrapolate human solutions and generate new ideas 42. Clustering algorithms can reduce a complex task’s solution space as humans explore possible solutions 43. Machines can unify similar solutions to mitigate statistical noise 44. Matching algorithms can match individuals and build efficient groups 45.
@@
This review focuses on integrating AI to bolster the collective intelligence of human groups, addressing a notable theoretical gap in understanding performance enhancements, particularly in hybrid human-AI configurations. With AI integration becoming pervasive across various sectors, exploring how this collaboration can unlock optimal capabilities is imperative. This exploration is essential for boosting productivity, fostering innovation, and ensuring that AI complements and enhances human skills rather than replacing them outright. The industry’s pivotal role in driving current AI-CI research underscores the absence of a comprehensive theoretical framework 46.
@@
In advancing the field of human-AI collective intelligence, prior research shows the necessity of formulating theories encompassing collective intelligence, combining both humans and AI 47. Developing such theoretical frameworks requires an in-depth comprehension and interpretation of human-AI systems, characterized by their high complexity and interrelated processes. Network science offers tools that enable us to understand the complexity of social systems 48 where a multitude of interacting elements give rise to the collective behavior of the whole system 49. To enhance our understanding of collective intelligence in the human-AI system and explore how “the whole is more than the sum of its parts” 50, 51, we integrate complexity science and network science approaches and propose a multilayer representation of the complex system involving human and AI agents.
Figure 1: A multilayer representation to untangle the processes in the complex system of human and AI agents, with three layers influencing each other: the cognition, physical, and information layers. External factors and the changing environment can influence the emergent collective intelligence of the whole system.
@@
Complex system thinking has been used to understand diverse biological, physical, and social domains 52, 53, 54. Inspired by this approach, a real-world collective intelligence system can be mapped into a multilayer network with three interconnected layers: cognition, physical, and information. In this network, nodes represent interacting agents, and links represent relationships between them. The construction of such networks and the meaning of the links are context-dependent. A node can also represent a group of agents instead of an individual agent in situations where the group behaves or is treated as a single entity within the network. Figure 1 illustrates a multiplex network 55, a special type of multilayer network where nodes remain the same but links differ across layers.
@@
The cognition layer, which contains mental processes, is only indirectly perceivable. Intelligent processes are involved during problem-solving, such as sense-making, remembrance, creation, coordination, and decision-making, often happening in the cognition layer. The links are hard to determine, but these processes exist and are fundamental to the emergence of collective intelligence in the whole system. In opposition to this indirectly perceivable layer is the physical layer, where humans and AI have tangible physical interactions. Information interaction in the information layer refers to exchanging information between agents through various communication channels. Besides intra-layer interactions, interactions in one layer can lead to information transfer and trigger interactions in other layers, represented as inter-layer links in the network 56. This interdependence and cross-layer influence are essential features of multilayer networks. Moreover, Even though our current framework revolves around pairwise interactions between nodes, recent advancements in the study of higher-order networks, going beyond pairwise interactions 57, may enrich our understanding of the collective emergence of intelligence.
@@
In addition to internal processes, it is essential to acknowledge the influence of the environment on the system 58. The human-AI complex system functions in a potentially dynamic environment. Previous research suggests that collective intelligent behavior depends on the environment 58, as certain behaviors or strategies that work well in one setting may not be as effective or relevant in another 59.
@@
Human-AI hybrid systems can be viewed as complex adaptive systems, continually evolving and adapting through interactions within dynamic environments 60. The concept of emergence in complex systems can be used to describe the phenomenon where new properties emerge at the collective level, which is not present at the level of individual components 61, 60, 62, 63. Here, the collective intelligence of the whole system can be seen as an emergent property superior to the micro-level individual intelligence. This property, which includes outcomes, abilities, characteristics, and behaviors, is aligned with and encompasses the major existing definitions of collective intelligence. It emerges through complex nonlinear relationships between the agents and is likely a result of bottom-up and top-down processes 16. The former encompasses the aggregation of group-member characteristics that foster collaboration. The latter includes structures, norms, and routines that govern collective behavior, thereby influencing the efficacy of coordination and collaboration 16. The modeling of the emergence of collective intelligence has leveraged analogies from a diverse array of fields 64, ranging from statistical physics 65 to neuroscience 66, both within and beyond.
Employing a multilayer representation of this complex system in a changing environment can facilitate a deeper understanding of the interactions and relationships between humans and AI agents, untangling the intricate processes involved in emerging and maintaining collective intelligence. Given the interdisciplinary nature of this approach and the variety of terminologies used, key terms and main concepts are listed in supplemental information Table S1 for more clarity. Table 2 provides a summary and breakdown of the components and features within the proposed multilayer network. The immediate benefit of this framework is that we can learn from the vast literature on multi-layer networks and the studies on their robustness67, adaptivity68, scalability69, resielence70, and interoperability71. We will discuss this further in the following sections.
@@
Task types can vary significantly, from generative activities such as brainstorming, which requires creative and divergent thinking, to analytical tasks such as solving Sudoku, which demands logical reasoning and pattern recognition 72. The diversity of the tasks requires different abilities, including memory utilization, creative imagination, sense-making, and critical analysis. In the following subsections, we focus on the details of group-member diversity’s influence, the structure and dynamics of interaction networks, and AI’s role in enhancing group performance.
@@
The relationship between group diversity and collective intelligence is highly complex due to diversity’s complexity and the range of its effects under different conditions. A group can exhibit surface-level diversities 73. The surface-level diversity comes from readily observable social categories, including gender, age, and ethnicity 74. The deep-level diversity refers to differences in psychological characteristics 75, including personality, cognitive thinking styles, and values.
Diversity benefits teams by enhancing creativity, improving decision-making, and expanding access to a broader talent pool.  76, 77. However, diversity can also have adverse effects, such as emotional conflict 78, 79, stress 80, poor work relationships, and poor overall performance 81. There is a lack of consensus on which aspects of group diversity are likely to result in positive outcomes and which aspects could produce negative outcomes 81, 82. The effects of diversity on group performance depend on the complexity of the task and various moderators, such as the density of the interaction network 83. A review underscores the need for in-depth research into surface and deep-level diversity within hybrid teams 82, exploring their impact across various outcomes and contexts over time.
@@
The evidence concerning the effect of gender diversity on team performance is equivocal and contingent upon various contextual factors 84. Past research found that gender-diverse teams outperform gender-homogeneous teams when perceived time pressure is low 85. One study showed gender diversity was negatively associated with performance, but only in large groups 74. Previous evidence strongly suggests that team collaboration is greatly improved by the presence of females in the group 84, 15. Gender-diverse teams produce more novel and higher-impact scientific ideas 86. Other studies show that gender composition seems to matter for team performance. However, when controlling for the individual abilities of team members, the relation between gender composition and team performance vanishes 87. One study points out that the “romance” of working together can benefit group performance 88. Gender-diverse groups perform better than homogeneous groups by decreasing relationship and task conflict 88.
@@
Age diversity may enhance group performance, but the benefits may depend on the context. Several studies have found that team age diversity positively correlated with team performance when completing complex tasks 89. Past research points out the positive effects were only seen under conditions of positive team climate and low age discrimination 89. Another study shows teams with more age diversity reported more age discrimination 90, associated with lower commitment and worse performance.
@@
There are inconsistent results in the case of ethnic diversity. One study 91 found racially diverse groups produced significantly more feasible and effective ideas than homogeneous groups. Another study found racial similarity in groups associated with higher self-rated productivity and commitment 92. Under the right conditions, teams may benefit from diversity in ethnicity and nationality 93.
@@
AI exhibits surface-level diversity through different anthropomorphic features in robots, such as voice, avatars, and human-like characteristics. Research has indicated that the gendering of AI can affect perceptions and trust levels in users. For instance, studies have found that users may exhibit more trust in AI personal assistants whose voice gender matches their own 94. Assigning a female gender to AI bots enhances their perceived humanness and acceptance 95. Previous studies have indicated a perception bias in AI agents based on gender: male AI agents tend to be viewed as more competent, whereas female AI agents are often perceived as warmer 96. The perceived gender of the machine can make the social dynamics in hybrid teams even more complicated. However, these gendered characteristics in AI also raise concerns about reinforcing stereotypes and biases 97. The influence of AI’s gendered traits on human interactions is an ongoing area of research, with implications for how AI is designed and utilized. Despite substantial research on AI gender, there is a notable gap in understanding the influence of other characteristics, such as AI age and ethnicity, on human interactions, highlighting the need for further exploration in this area.
@@
Some work proposes that investigations into the relationship between personality and work-related behaviors should expand beyond the linearity assumptions 98, showing that extraversion, agreeableness, and conscientiousness have inverted U-shaped relationships with peer-rated contributions to teamwork 98. Similar work demonstrates a curvilinear relationship between a team’s average proactive personality and performance 99. This body of work also identifies the moderating effects of personality dispersion, team potency, and cohesion on this relationship 99. Previous research has found that a team’s collective openness to experience and emotional stability moderates how task conflict affects team performance 100. Specifically, in teams with high openness or emotional stability, task conflict positively influences performance 100.
@@
Team cognitive style diversity refers to the variation in team members’ ways of encoding, organizing, and processing information 101. Research indicates its significance in fostering innovation within teams 102. A certain level of cognitive diversity contributes to collective intelligence by providing diverse cognitive inputs and viewpoints necessary for task work 102. However, excessive diversity may lead to high coordination costs 103, as team members with differing perspectives struggle to understand each other 104, 105. Studies suggest that the relationship between cognitive style diversity and collective intelligence is curvilinear, forming an inverted U-shape 101. Additionally, the beneficial effects of cognitive diversity might be mediated through factors such as task reflexivity and relationship conflict 106.
@@
Values are internalized beliefs that can guide behavior and enhance motivation 107. Past research found higher diversity in values was associated with more conflict 108 in relationships, tasks, and processes. Overall, existing research consistently indicates that value congruence within teams positively influences performance, satisfaction, and conflict 82 and moderates the effects of informational diversity 109.
@@
While AI systems can display various behaviors or responses based on their programming and training, this diversity is a product of different algorithms, models, and data inputs. Unlike humans, AI does not have personal experiences, innate personality traits, cognitive thinking, or beliefs contributing to deep-level diversity. Although AI lacks inherent deep-level traits, AI might influence deep-level diversity in human teams, primarily by shaping social interactions and decision-making processes. The optimal composition of teams remains uncertain 110.
@@
Interaction distinguishes a group from a mere collection of individuals 58: one person’s behavior forms the basis for the responses of others 111. Therefore, the interaction between two members establishes a link in the interaction network. In terms of the group’s interaction network structure, past research suggests that the small-worldness of a collaboration network (small network diameter and high clustering) improves performance up to a certain point, after which the effect will reverse 112. An “inverted U” relationship is found between connectedness and performance 113, suggesting an optimum number of connections for any given size of the collaboration network. A modular structure (a large number of connections within small groups that are loosely connected) can increase efficiency 114. A study on forbidden triads among jazz groups shows that heavily sided open triangles are associated with lower success (measured by the number of group releases) 115.
AI can contribute to human groups in various ways by augmenting existing human skills or complementing capabilities that humans lack. AI has attributes beyond humans, such as more extended memory, higher computational speed, and a more vital ability to work with large, diverse datasets. AI’s augmentation of human cognition enables teams to navigate complex situations effectively 116. AI may also help reduce the harm of implicit bias in humans 117 and shape better decisions. Taking advantage of the differences between humans and AI can contribute to the performance of hybrid human-AI groups. In the following sections, we discuss the rules and incentives that govern the interaction processes and the structure and dynamics of the group’s interaction network.
@@
Decision rules function as prescribed norms that direct interactions and play a crucial role in shaping the communication and information integration within a group 118. Previous research has focused on disentangling decision rules guiding the team’s interaction, ultimately fostering synergy. Some literature emphasizes that the decision rules themselves are intelligent 119. In addition to the rules guiding behavior, the group members interact with each other under specific incentive schemes. The intrinsic and extrinsic incentives are the driving forces 120 for the agents to behave in a direction such that the social network becomes dynamic and moves towards a collective goal. Some research proposes that an incentive scheme rewarding accurate predictions by a minority can foster optimal diversity and collective predictive accuracy 121.
AI can augment human cognition 122 to help teams adapt to complexity. While AI can simulate certain human-like incentives through programming and learning algorithms 123, fully capturing the depth and nuance of human behavior, such as irrationality or altruism, is challenging. A recent study applying a Turing test to compare AI chatbots and humans found that AI chatbots exhibited more altruistic and cooperative behavior than their human counterparts in behavioral games 124. Machines can be programmed to mimic human behaviors, but whether they genuinely “exhibit” them like humans is debatable. In collective intelligence, integrating such human-like behaviors in AI may enhance collaboration and decision-making within diverse teams. However, the extent to which AI can be truly self-aware and authentically replicate complex human traits remains a subject of ongoing research and philosophical debate.

### v4

@@
Hao Cui School of Sociology, University College Dublin, Dublin, Ireland Geary Institute for Public Policy, University College Dublin, Dublin, Ireland Taha Yasseri School of Sociology, University College Dublin, Dublin, Ireland Geary Institute for Public Policy, University College Dublin, Dublin, Ireland School of Social Sciences and Philosophy, Trinity College Dublin, Dublin, Ireland
@@
Current societal challenges exceed the capacity of humans operating either alone or collectively. As AI evolves, its role within human collectives will vary from an assistive tool to a participatory member. Humans and AI possess complementary capabilities that, together, can surpass the collective intelligence of either humans or AI in isolation. However, the interactions in human-AI systems are inherently complex, involving intricate processes and interdependencies. This review incorporates perspectives from complex network science to conceptualize a multilayer representation of human-AI collective intelligence, comprising cognition, physical, and information layers. Within this multilayer network, humans and AI agents exhibit varying characteristics; humans differ in diversity from surface-level to deep-level attributes, while AI agents range in degrees of functionality and anthropomorphism. We explore how agents’ diversity and interactions influence the system’s collective intelligence and analyze real-world instances of AI-enhanced collective intelligence. We conclude by considering potential challenges and future developments in this field.
@@ Introduction
The definition of artificial intelligence has undergone numerous evolutions in the past semi-century. Artificial intelligence is the simulation of human intelligence processes by machines 23, especially computer systems. Previous research considers the classification of AI in distinct ways 24. The first classification categorizes AI based on its human-like cognitive abilities, such as thinking and feeling, and contains four primary AI types 24: reactive AI, limited memory AI, theory of mind AI, and self-aware AI. Limited-memory AI is the most common at the current time, ranging from virtual assistants to chatbots to self-driving vehicles. Such AI can learn from historical data to recognize patterns, generate new knowledge and understanding, and inform subsequent decisions.
Yet, another classification is a technology-oriented approach that categorizes AI into Artificial Narrow Intelligence (ANI), Artificial General Intelligence (AGI), and Artificial Superintelligence (ASI). ANI involves creating computer systems capable of performing specific tasks like human intelligence 25, but often surpass humans in efficiency and accuracy 26. ANI machines have a narrow range of capabilities and can represent all existing AI. AGI refers to machines that exhibit the human ability to learn, perceive, and understand a wide range of intellectual tasks 27, 28. Finally, ASI’s primary goal is to develop a machine with cognitive abilities higher than those of humans 24.
Existing AI can perform tasks including but not limited to complex calculations, language translation, facial recognition, and financial market prediction. Recent developments in generative AI, such as OpenAI’s text-to-video model Sora (openai.com/sora), exemplify AI’s potential in creative industries. As AI technology evolves, it is poised to bring groundbreaking advancements across various fields. Beyond content creation, as AI deepens the understanding of the physical world and develops its simulation ability at the perceptual and cognitive level, it is possible to develop more super-intelligent tools in various fields.
Current AI systems can process vast amounts of data on a scale far beyond human capabilities. However, many real-world challenges cannot yet be solved solely by AI. Currently, AI lacks the deep conceptual and emotional understanding humans possess about objects and experiences 29. AI cannot interpret human language’s nuances, including contextual, symbolic, and cultural meanings 30, 31. AI cannot be an ethical decision-maker because it lacks the human attributes of intentionality, care, and responsibility 29. Therefore, combining human insight with AI’s analytical power is crucial for addressing complex real-world challenges, leveraging the strengths to compensate for each other’s weaknesses.
@@
Even though the field of collective intelligence initially focused solely on groups of people, in recent years, it has gradually expanded to include AI as group members in a new framework referred to as “hybrid intelligence” 32. Proponents of the hybrid intelligence perspective stress that humans and AI can connect in ways that allow them to collectively act more intelligently and achieve unreachable goals by any individual entities alone 33. Researchers investigating this hybrid collective intelligence explore “how people and computers can be connected so that collectively they act more intelligently than any person, group, or computer has ever done before” 34.
Acknowledging the complementary capabilities of humans and AI as discussed in the previous section, researchers identify the need for developing socio-technological ensembles of humans and intelligent machines that possess the ability to collectively achieve superior results and continuously improve by learning from each other 35. Previous research provides evidence that teaming humans with AI has the potential to achieve complementary team performance (CTP), a level of performance that AI or humans cannot reach in isolation 36. For instance, a study demonstrates that humans can use contextual information to adjust the AI’s decision 37. Research on mixed teams composed of humans and AI shows hybrid teams could achieve higher performance metrics, such as team situational awareness and score, than all-human teams 36. Finally, recent advancements in developing Large Language Models (LLMs) have also been proposed to reshape collective intelligence 38.
@@
For example, machine learning and automation can increase the efficiency and scalability of collective intelligence in Citizen Science projects, where citizens volunteer to help scientists tag and classify their large-scale datasets 39. Members of a crowd with different interests and areas of competence can be matched to other tasks by recommendation systems 40. Generative algorithms can extrapolate human solutions and generate new ideas 41. Clustering algorithms can reduce a complex task’s solution space as humans explore possible solutions 42. Machines can unify similar solutions to mitigate statistical noise 43. Matching algorithms can match individuals and build efficient groups 44.
@@
This review focuses on integrating AI to bolster the collective intelligence of human groups, addressing a notable theoretical gap in understanding performance enhancements, particularly in hybrid human-AI configurations. With AI integration becoming pervasive across various sectors, exploring how this collaboration can unlock optimal capabilities is imperative. This exploration is essential for boosting productivity, fostering innovation, and ensuring that AI complements and enhances human skills rather than replacing them outright. The industry’s pivotal role in driving current AI-CI research underscores the absence of a comprehensive theoretical framework 45.
@@
In advancing the field of human-AI collective intelligence, prior research shows the necessity of formulating theories encompassing collective intelligence, combining both humans and AI 46. Developing such theoretical frameworks requires an in-depth comprehension and interpretation of human-AI systems, characterized by their high complexity and interrelated processes. Network science offers tools that enable us to understand the complexity of social systems 47 where a multitude of interacting elements give rise to the collective behavior of the whole system 48. To enhance our understanding of collective intelligence in the human-AI system and explore how “the whole is more than the sum of its parts” 49, 50, we integrate complexity science and network science approaches and propose a multilayer representation of the complex system involving human and AI agents.
@@
Complex system thinking has been used to understand diverse biological, physical, and social domains 51, 52, 53. Inspired by this approach, a real-world collective intelligence system can be mapped into a multilayer network with three interconnected layers: cognition, physical, and information. In this network, nodes represent interacting agents, and links represent relationships between them. The construction of such networks and the meaning of the links are context-dependent. A node can also represent a group of agents instead of an individual agent in situations where the group behaves or is treated as a single entity within the network. Figure 1 illustrates a multiplex network 54, a special type of multilayer network where nodes remain the same but links differ across layers.
@@
The cognition layer, which contains mental processes, is only indirectly perceivable. Intelligent processes are involved during problem-solving, such as sense-making, remembrance, creation, coordination, and decision-making, often happening in the cognition layer. The links are hard to determine, but these processes exist and are fundamental to the emergence of collective intelligence in the whole system. In opposition to this indirectly perceivable layer is the physical layer, where humans and AI have tangible physical interactions. Information interaction in the information layer refers to exchanging information between agents through various communication channels. Besides intra-layer interactions, interactions in one layer can lead to information transfer and trigger interactions in other layers, represented as inter-layer links in the network 55. This interdependence and cross-layer influence are essential features of multilayer networks. Moreover, Even though our current framework revolves around pairwise interactions between nodes, recent advancements in the study of higher-order networks, going beyond pairwise interactions 56, may enrich our understanding of the collective emergence of intelligence.
@@
In addition to internal processes, it is essential to acknowledge the influence of the environment on the system 57. The human-AI complex system functions in a potentially dynamic environment. Previous research suggests that collective intelligent behavior depends on the environment 57, as certain behaviors or strategies that work well in one setting may not be as effective or relevant in another 58.
@@
Human-AI hybrid systems can be viewed as complex adaptive systems, continually evolving and adapting through interactions within dynamic environments 59. The concept of emergence in complex systems can be used to describe the phenomenon where new properties emerge at the collective level, which is not present at the level of individual components 60, 59, 61, 62. Here, the collective intelligence of the whole system can be seen as an emergent property superior to the micro-level individual intelligence. This property, which includes outcomes, abilities, characteristics, and behaviors, is aligned with and encompasses the major existing definitions of collective intelligence. It emerges through complex nonlinear relationships between the agents and is likely a result of bottom-up and top-down processes 16. The former encompasses the aggregation of group-member characteristics that foster collaboration. The latter includes structures, norms, and routines that govern collective behavior, thereby influencing the efficacy of coordination and collaboration 16. The modeling of the emergence of collective intelligence has leveraged analogies from a diverse array of fields 63, ranging from statistical physics 64 to neuroscience 65, both within and beyond.
Employing a multilayer representation of this complex system in a changing environment can facilitate a deeper understanding of the interactions and relationships between humans and AI agents, untangling the intricate processes involved in emerging and maintaining collective intelligence. Given the interdisciplinary nature of this approach and the variety of terminologies used, key terms and main concepts are listed in supplemental information Table S1 for more clarity. Table 2 provides a summary and breakdown of the components and features within the proposed multilayer network. The immediate benefit of this framework is that we can learn from the vast literature on multilayer networks and the studies on their robustness66, adaptivity67, scalability68, resilience69, and interoperability70. We will discuss this further in the following sections.
@@
Task types can vary significantly, from generative activities such as brainstorming, which requires creative and divergent thinking, to analytical tasks such as solving Sudoku, which demands logical reasoning and pattern recognition 71. The diversity of the tasks requires different abilities, including memory utilization, creative imagination, sense-making, and critical analysis. In the following subsections, we focus on the details of group-member diversity’s influence, the structure and dynamics of interaction networks, and AI’s role in enhancing group performance.
@@
The relationship between group diversity and collective intelligence is highly complex due to diversity’s complexity and the range of its effects under different conditions. A group can exhibit surface-level diversities 72. The surface-level diversity comes from readily observable social categories, including gender, age, and ethnicity 73. The deep-level diversity refers to differences in psychological characteristics 74, including personality, cognitive thinking styles, and values.
Diversity benefits teams by enhancing creativity, improving decision-making, and expanding access to a broader talent pool.  75, 76. However, diversity can also have adverse effects, such as emotional conflict 77, 78, stress 79, poor work relationships, and poor overall performance 80. There is a lack of consensus on which aspects of group diversity are likely to result in positive outcomes and which aspects could produce negative outcomes 80, 81. The effects of diversity on group performance depend on the complexity of the task and various moderators, such as the density of the interaction network 82. A review underscores the need for in-depth research into surface and deep-level diversity within hybrid teams 81, exploring their impact across various outcomes and contexts over time.
@@
The evidence concerning the effect of gender diversity on team performance is equivocal and contingent upon various contextual factors 83. Past research found that gender-diverse teams outperform gender-homogeneous teams when perceived time pressure is low 84. One study showed gender diversity was negatively associated with performance, but only in large groups 73. Previous evidence strongly suggests that team collaboration is greatly improved by the presence of females in the group 83, 15. Gender-diverse teams produce more novel and higher-impact scientific ideas 85. Other studies show that gender composition seems to matter for team performance. However, when controlling for the individual abilities of team members, the relation between gender composition and team performance vanishes 86. One study points out that the “romance” of working together can benefit group performance 87. Gender-diverse groups perform better than homogeneous groups by decreasing relationship and task conflict 87.
@@
Age diversity may enhance group performance, but the benefits may depend on the context. Several studies have found that team age diversity positively correlated with team performance when completing complex tasks 88. Past research points out the positive effects were only seen under conditions of positive team climate and low age discrimination 88. Another study shows teams with more age diversity reported more age discrimination 89, associated with lower commitment and worse performance.
@@
There are inconsistent results in the case of ethnic diversity. One study 90 found racially diverse groups produced significantly more feasible and effective ideas than homogeneous groups. Another study found racial similarity in groups associated with higher self-rated productivity and commitment 91. Under the right conditions, teams may benefit from diversity in ethnicity and nationality 92.
@@
AI exhibits surface-level diversity through different anthropomorphic features in robots, such as voice, avatars, and human-like characteristics. Research has indicated that the gendering of AI can affect perceptions and trust levels in users. For instance, studies have found that users may exhibit more trust in AI personal assistants whose voice gender matches their own 93. Assigning a female gender to AI bots enhances their perceived humanness and acceptance 94. Previous studies have indicated a perception bias in AI agents based on gender: male AI agents tend to be viewed as more competent, whereas female AI agents are often perceived as warmer 95. The perceived gender of the machine can make the social dynamics in hybrid teams even more complicated. However, these gendered characteristics in AI also raise concerns about reinforcing stereotypes and biases 96. The influence of AI’s gendered traits on human interactions is an ongoing area of research, with implications for how AI is designed and utilized. Despite substantial research on AI gender, there is a notable gap in understanding the influence of other characteristics, such as AI age and ethnicity, on human interactions, highlighting the need for further exploration in this area.
@@
Some work proposes that investigations into the relationship between personality and work-related behaviors should expand beyond the linearity assumptions 97, showing that extraversion, agreeableness, and conscientiousness have inverted U-shaped relationships with peer-rated contributions to teamwork 97. Similar work demonstrates a curvilinear relationship between a team’s average proactive personality and performance 98. This body of work also identifies the moderating effects of personality dispersion, team potency, and cohesion on this relationship 98. Previous research has found that a team’s collective openness to experience and emotional stability moderates how task conflict affects team performance 99. Specifically, in teams with high openness or emotional stability, task conflict positively influences performance 99.
@@
Team cognitive style diversity refers to the variation in team members’ ways of encoding, organizing, and processing information 100. Research indicates its significance in fostering innovation within teams 101. A certain level of cognitive diversity contributes to collective intelligence by providing diverse cognitive inputs and viewpoints necessary for task work 101. However, excessive diversity may lead to high coordination costs 102, as team members with differing perspectives struggle to understand each other 103, 104. Studies suggest that the relationship between cognitive style diversity and collective intelligence is curvilinear, forming an inverted U-shape 100. Additionally, the beneficial effects of cognitive diversity might be mediated through factors such as task reflexivity and relationship conflict 105.
@@
Values are internalized beliefs that can guide behavior and enhance motivation 106. Past research found higher diversity in values was associated with more conflict 107 in relationships, tasks, and processes. Overall, existing research consistently indicates that value congruence within teams positively influences performance, satisfaction, and conflict 81 and moderates the effects of informational diversity 108.
@@
While AI systems can display various behaviors or responses based on their programming and training, this diversity is a product of different algorithms, models, and data inputs. Unlike humans, AI does not have personal experiences, innate personality traits, cognitive thinking, or beliefs contributing to deep-level diversity. Although AI lacks inherent deep-level traits, AI might influence deep-level diversity in human teams, primarily by shaping social interactions and decision-making processes. The optimal composition of teams remains uncertain 109.
@@
Interaction distinguishes a group from a mere collection of individuals 57: one person’s behavior forms the basis for the responses of others 110. Therefore, the interaction between two members establishes a link in the interaction network. In terms of the group’s interaction network structure, past research suggests that the small-worldness of a collaboration network (small network diameter and high clustering) improves performance up to a certain point, after which the effect will reverse 111. An “inverted U” relationship is found between connectedness and performance 112, suggesting an optimum number of connections for any given size of the collaboration network. A modular structure (a large number of connections within small groups that are loosely connected) can increase efficiency 113. A study on forbidden triads among jazz groups shows that heavily sided open triangles are associated with lower success (measured by the number of group releases) 114.
AI can contribute to human groups in various ways by augmenting existing human skills or complementing capabilities that humans lack. AI has attributes beyond humans, such as more extended memory, higher computational speed, and a more vital ability to work with large, diverse datasets. AI’s augmentation of human cognition enables teams to navigate complex situations effectively 19. AI may also help reduce the harm of implicit bias in humans 115 and shape better decisions. Taking advantage of the differences between humans and AI can contribute to the performance of hybrid human-AI groups. In the following sections, we discuss the rules and incentives that govern the interaction processes and the structure and dynamics of the group’s interaction network.
@@
Decision rules function as prescribed norms that direct interactions and play a crucial role in shaping the communication and information integration within a group 116. Previous research has focused on disentangling decision rules guiding the team’s interaction, ultimately fostering synergy. Some literature emphasizes that the decision rules themselves are intelligent 117. In addition to the rules guiding behavior, the group members interact with each other under specific incentive schemes. The intrinsic and extrinsic incentives are the driving forces 118 for the agents to behave in a direction such that the social network becomes dynamic and moves towards a collective goal. Some research proposes that an incentive scheme rewarding accurate predictions by a minority can foster optimal diversity and collective predictive accuracy 119.
AI can augment human cognition 120 to help teams adapt to complexity. While AI can simulate certain human-like incentives through programming and learning algorithms 121, fully capturing the depth and nuance of human behavior, such as irrationality or altruism, is challenging. A recent study applying a Turing test to compare AI chatbots and humans found that AI chatbots exhibited more altruistic and cooperative behavior than their human counterparts in behavioral games 122. Machines can be programmed to mimic human behaviors, but whether they genuinely “exhibit” them like humans is debatable. In collective intelligence, integrating such human-like behaviors in AI may enhance collaboration and decision-making within diverse teams. However, the extent to which AI can be truly self-aware and authentically replicate complex human traits remains a subject of ongoing research and philosophical debate.